# Learning Stable and Passive Neural Differential Equations

### v1

@@ I Introduction
Given an Lyapunov function $V(x)$, one needs to construct a vector field $f(x)$ such that the Lyapunov inequality $\nabla^{\top}V(x)f(x)<0$ holds for nonzero $x$. An intuitive approach is the gradient flow, i.e., $f(x)=-\alpha(x)\nabla V(x)$ with $\alpha(x)>0$. However, it is often restrictive as there are infinitely many of descent directions. In [15], a projection method is applied to resolve the case where Lyapunov inequality is infeasible. There are two main drawbacks: 1) the vector field becomes non-smooth due to projection; 2) when projection is activated, the vector field is reduced to a gradient flow of a convex function.
To address the above issues, we take a recently developed bi-Lipschitz network from [18] to construct a Lyapunov function $H(x)=0.5|g(x)|^{2}$. Since $g$ is bi-Lipschitz, then it is also invertible, implying that the level set of $H$ is homeomorphic to a unit ball. Unlike the ICNN-based Lyapunov function, the level sets of $H$ can be non-convex. Thanks to the bi-Lipschitz property of $g$, both quadratic lower and upper bounds are guaranteed automatically. We then construct the vector field as $f(x)=[J(x)-R(x)]\nabla H(x)$ where $J$ is skew symmetric and $R$ is positive semidefinite. The above $f$ is a smooth field of the descent vectors of $H$. By parameterizing $J$ and $R$ using neural networks, one can learn flexible stable dynamics. Moreover, this structure can also be realized as the Hamiltonian form that widely exists in physical systems [19, 20].
@@
  • It has certificated global Lyapunov stability and exponential stability, with respect to an equilibrium point. The equilibrium point can either be imposed as a prior knowledge or learned by the data.
@@
Notation. Given a $x\in\mathbb{R}^{n}$, we denote a ball around $x$ with radium $\epsilon$ by $\mathcal{B}_{\epsilon}(x):=\{y\in\mathbb{R}^{n}\mid|y-x|\leq\epsilon\}$ where ||$\cdot$|| ⋅ | is the Euclidean norm. We define the gradient transpose operator as $\nabla_{x}:=(\partial/\partial x)^{\top}$. The subscript will be omitted when it is clear from the context.
@@ II-A Problem setup
We also consider learning $v=f(x,u)$ and $y=h(x)$ from a dataset $\mathcal{D}=\{(x_{i},v_{i},u_{i},y_{i})\mid x_{i},v_{i}\in\mathbb{R}^{n},u_{i},y_{i}\in\mathbb{R}^{m}\}$ such that the following controlled system
@@ II-B Preliminaries
These features make the PLNet a promising candidate for Lyapunov functions.
@@ III-A Stable Hamiltonian models
$|x(t)-x^{\star}|\leq\frac{\sqrt{2H(x(t))}}{\nu}\leq\frac{\sqrt{2H(x(0))}}{\nu}\leq\frac{\nu}{\mu}|x(0)-x^{\star}|,$
@@
We now parameterize the matrix functions $J(x)$ and $R(x)$. Let $S(x)$ and $L(x)$ be two learnable multi-layer perceptions (MLPs), which are mappings from $R^{n}$ to $\mathbb{R}^{n\times n}$. We then construct
@@ IV Experiments
$\ell(f)=\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\bigl{|}v_{i}-f(x_{i})\bigr{|}^{2}.$ (19)
with $|\mathcal{D}|$ as the number of elements in $\mathcal{D}$, where $\mathcal{D}$ is either $\mathcal{D}_{\mathrm{train}}$ or $\mathcal{D}_{\mathrm{test}}$.
@@ IV-A Model architecture and training details
where $\texttt{ReLU}(x)=\max(0,x)$. If the vector filed $\hat{f}$ satisfies the following Lyapunov inequality
$\nabla V(x)^{\top}\hat{f}(x)\leq-\alpha V(x),$ (21)
we have $f(x)=\hat{f}(x)$. For the region where the above Lyapunov inequality does not hold, the vector field $f$ then takes the descent direction of $V$, i.e.,
$f(x)=-\alpha\frac{V(x)}{|\nabla V(x)|^{2}}\nabla V(x).$ (22)
For the vector field $\hat{f}$, we choose a 4-100-100-4 multi-layer perception (MLP), i.e., a network with input and output dimension of 4 and two hidden layer with each has 100 neurons. For the Lyapunov function $V$, we consider two the following two cases extracted from [15]:
@@
For the proposed SHND model, we use the bi-Lipschitz network (15) from [18] with architecture 4-32-32-4 and Lipschitz bound $(\mu,\nu)=(0.1,2)$ to construct the Hamiltonian $H(x)$. For the matrix functions $J(x)$ and $R(x)$, we first take a 4-90-90-32 MLP, then partition the output into two $4\times 4$ matrices $L$ and $S$, and finally compute $J,R$ via (16) with $\epsilon=0.01$. As shown in Thm. 1, the SHND model can provide global exponential stability guarantee for the equilibrium $x^{\star}$. We will illustrate this point in the next section. For a fair comparison, our SHND model has a similar amount of learnable model parameters (around 15K) as the SD-MLP and SD-ICNN.
@@ IV-B Results and discussions
We can observed that our model has much smaller simulation errors compared to SD-ICNN and SD-MLP. In particular, the simulation error of SHND converges to 0 as the simulation time increases, demonstrating that our model can provide stability guarantee for the known equilibrium $x^{\star}$. For SD-ICNN model, as shown in Fig. 5, with slight changes in the initial condition, it converges to a different equilibrium. The SD-MLP produces oscillating behaviors as its Lyapunov function is not positive definite. Also, its simulation error is much larger than the other two approaches.
@@
Figure 5: Simulation trajectories of different models with two initial states. The SD-MLP model generates oscialtting behaviors. Both SD-ICNN and the proposed SHND converge. But the SD-ICNN converges to a non-zero equilibrium point despite a small change in the initial condition.
@@ V Conclusion
In this paper, we have proposed a new class of neural differential equations called stable Hamiltonian neural dynamics (SHND), which has built-in stability with respect to an equilibrium. Empirical results on a vector field fitting problem demonstrate its effectiveness. For future directions, we will consider to learn stable models from time-series data and also explore applications in learning passivity-based controllers.

### v2

@@ I Introduction
Given an Lyapunov function $V(x)$, one needs to construct a vector field $f(x)$ such that the Lyapunov inequality $\nabla^{\top}V(x)f(x)<0$ holds for nonzero $x$. An intuitive approach is the gradient flow, i.e., $f(x)=-\alpha(x)\nabla V(x)$ with $\alpha(x)>0$. However, it is often restrictive as there are infinitely many of descent directions. In [15], a projection method is applied to resolve the case where Lyapunov inequality is infeasible.
In this work, we take a recently developed bi-Lipschitz network $g$ from [18] to construct a Lyapunov function $H(x)=0.5|g(x)|^{2}$. Since $g$ is bi-Lipschitz, then it is also invertible, implying that the level set of $H$ is homeomorphic to a unit ball. Unlike the ICNN-based Lyapunov function, the level sets of $H$ can be non-convex. Thanks to the bi-Lipschitz property of $g$, both quadratic lower and upper bounds are guaranteed automatically. We then construct the vector field as $f(x)=[J(x)-R(x)]\nabla H(x)$ where $J$ is skew symmetric and $R$ is positive definite. The above $f$ is a smooth field of the descent vectors of $H$. By parameterizing $J$ and $R$ using neural networks, one can learn flexible stable dynamics. Moreover, this structure can also be realized as the Hamiltonian form that widely exists in physical systems [19, 20].
@@
  • It has certified global Lyapunov stability and exponential stability by the parameterized Lyapunov function, with respect to an equilibrium point. The equilibrium point can either be imposed as a prior knowledge or learned by the data.
@@
Notation. Given a $x\in\mathbb{R}^{n}$, we denote a ball around $x$ with radium $\epsilon$ by $\mathcal{B}_{\epsilon}(x):=\{y\in\mathbb{R}^{n}\mid|y-x|\leq\epsilon\}$ where ||$\cdot$|| ⋅ | is the Euclidean norm. We denote $A\succeq B$ if and only if the matrix $(A-B)$ is positive semi-definite. We define the gradient transpose operator as $\nabla_{x}:=(\partial/\partial x)^{\top}$. The subscript will be omitted when it is clear from the context.
@@ II-A Problem setup
We also consider learning $v=f(x,u)$ and $y=h(x)$ from a dataset $\mathcal{D}=\{(x_{i},v_{i},u_{i},y_{i})\mid x_{i},v_{i}\in\mathbb{R}^{n},u_{i},y_{i}\in\mathbb{R}^{m}\}$ such that the following controlled system in the input-state-output form
@@ II-B Preliminaries
These features make the PLNet a promising candidate for learning Lyapunov functions, e.g. using the methods of [28, 17]. In this paper we focus on learning stable dynamics by parameterising flows that are descent directions for such a Lyapunov function, and consider further extensions to passive dynamics.
@@ III-A Stable Hamiltonian models
$|x(t)-x^{\star}|\leq\frac{\sqrt{2H(x(t))}}{\mu}\leq\frac{\sqrt{2H(x(0))}}{\mu}\leq\frac{\nu}{\mu}|x(0)-x^{\star}|,$
@@
We now parameterize the matrix functions $J(x)$ and $R(x)$. Let $S(x)$ and $L(x)$ be two learnable multi-layer perceptions (MLPs), which are mappings from $\mathbb{R}^{n}$ to $\mathbb{R}^{n\times n}$. We then construct
@@ IV Experiments
$\ell(f)=\frac{1}{n_{\mathcal{D}}}\sum_{i=1}^{n_{\mathcal{D}}}\bigl{|}v_{i}-f(x_{i})\bigr{|}^{2}.$ (19)
with $n_{\mathcal{D}}$ as the number of elements in $\mathcal{D}$, where $\mathcal{D}$ is either $\mathcal{D}_{\mathrm{train}}$ or $\mathcal{D}_{\mathrm{test}}$.
@@ IV-A Model architecture and training details
where $\texttt{ReLU}(x)=\max(0,x)$. For the vector field $\hat{f}$, we choose a 4-100-100-4 multi-layer perception (MLP), i.e., a network with input and output dimension of 4 and two hidden layer with each has 100 neurons. For the Lyapunov function $V$, we consider the following two cases extracted from [15]:
@@
For the proposed SHND model, we use the bi-Lipschitz network (15) from [18] with architecture 4-32-32-4 and Lipschitz bound $(\mu,\nu)=(0.1,2)$ to construct the Hamiltonian $H(x)$. For the matrix functions $J(x)$ and $R(x)$, we first take a 4-90-90-32 MLP, then partition the output into two $4\times 4$ matrices $L$ and $S$, and finally compute $J,R$ via (16) with $\epsilon=0.01$. As shown in Thm 1, the SHND model can provide global exponential stability guarantee for the equilibrium $x^{\star}$. We will illustrate this point in the next section. For a fair comparison, our SHND model has a similar amount of learnable model parameters (around 15K) as the SD-MLP and SD-ICNN.
@@ IV-B Results and discussions
We can observe that our model has much smaller simulation errors compared to SD-ICNN and SD-MLP. In particular, the simulation error of SHND converges to 0 as the simulation time increases, demonstrating that our model can provide stability guarantee for the known equilibrium $x^{\star}$. For SD-ICNN model, as shown in Fig. 5, with slight changes in the initial condition, it converges to a different equilibrium. The SD-MLP produces oscillating behaviors as its Lyapunov function is not positive definite. Also, its simulation error is much larger than the other two approaches.
@@
Figure 5: Simulation trajectories of different models with two initial states. The SD-MLP model generates oscillating behaviors. Both SD-ICNN and the proposed SHND converge. But the SD-ICNN converges to a non-zero equilibrium point despite a small change in the initial condition.
@@ V Conclusion
In this paper, we have proposed a new class of neural differential equations called stable Hamiltonian neural dynamics (SHND), which has built-in stability w.r.t. an equilibrium. Empirical results on a vector field fitting problem demonstrate its effectiveness. For future directions, we will consider to learn stable models from time-series data and also explore applications in learning passivity-based controllers.